Which graph represents the parametric equations 

x = t – 4

y = t2 + 5

–1 ≤ t ≤ 2


1
Expert's answer
2021-03-29T08:55:24-0400

Given:

"x=t-4\\y=t^2+5"


So to find equation of the curve we start with y,

and using a basic conversion in the given equation,

"y=t^2+5\\y=t^2+5-16+16
\\y=(t^2-16)+21
\\y=(t-4)(t+4)+21
\\y=(t-4)(t-4+8)+21"


Now putting,

"x=t-4"

we further get,

"y=x(x+4)+21
\\y=x^2+8x+21" .............is the required curve. Now plotting on the graph


–1 ≤ t ≤ 2

we get,



This graph represents the given parametric equations.